OpenStudy (anonymous):

Easy Medal!!!!!!!! What is the area, in square meters, of a right triangle with a hypotenuse of 65 meters and a leg of 25 meters? a. 1,500 m2 c. 1,950 m2 b. 8,125 m2 d. 750 m2

OpenStudy (anonymous):

I assume lrg is base area= 1/2 * base * height 1/2 * 25 * h h= sqrt 65^2-25^2 h=sqrt 3600 h=60 Area therefore = 1/2 * 25 * 60 750m^2

OpenStudy (anonymous):

Area = (1/2) base x height have base and hypotenuse, get height from base^2 + height^2 = hypotenuse^2